OpenStudy (anonymous):

distributive property and combining like terms 4x-8+x=2

OpenStudy (across):

We have the equation: \[4x-8+x=2\]We combine like terms: \[5x-8=2\]We isolate x: \[5x=10\]We simplify x: \[x=2\]EASY

OpenStudy (anonymous):

Okay, so we have: 4x - 8 + x = 2 We set the equation to zero. 4x - 8 + x - 2 = 0 Now if we add like terms, we have: 5x - 10 = 0 We can remove a 5 from both terms since they have 5 in them. This gives us: 5(x-2) = 0 That is the distributive property. By simple algebra 5x - 10 = 0 becomes x = 2 Hope this helps clarify things. http://www.tutorsean.net